Let A = {1, 2, 3, 4} and B = {0, 3, 6, 8, 12, 15}. Consider a rule f (x) = x² - 1, x∈A, then show that f is a mapping from A to B.



1
Expert's answer
2022-06-13T23:23:49-0400

Let's consider a rule "f (x) = x\u00b2 - 1, x\u2208A" and find an image of every element in A.

At "x=1:" "f(1)=1^2-1=0\in B."

At "x=2: \: f(2)=2^2-1=4-1=3\in B."

At "x=3: \: f(3)=3^2-1=9-1=8\in B."

At "x=4: \: f(4)=4^2-1=16-1=15\in B."


We see that every element "x\u2208A" has an image "f(x)\u2208B" and the image is unique.

It shows that "f" is a mapping from A to B.